OpenStudy (zuha):

Find each unit price. Which is the better buy? Explain. Electrical wire: $1.49 for 3 ft. $.69 for 18 in

OpenStudy (zuha):

please help !!

OpenStudy (zuha):

I'm waiting!!

OpenStudy (texaschic101):

1.49/3 = x/1 cross multiply 3x = 1.49 x = 1.49/3 x = 0.496 (rounds to 50 cents per foot) .69/18 = x/1 cross multiply 18x = .69 x = .69/18 x = .038 per inch 12 inches in a foot x = .038(12) = 0.456 (rounds to 46 cents per foot) better deal is the 2nd one

OpenStudy (zuha):

thanks

OpenStudy (texaschic101):

no problem :)